OpenStudy (anonymous):

What is the y-coordinate of the solution of the system? { x + y = 3 { 4x - y = -38

OpenStudy (fortytherapper):

Let's first rearrange each equation to where it says y = ______

OpenStudy (anonymous):

Alright i have no clue what y = lol

OpenStudy (mathstudent55):

Because both equations are written in the same form, and they line up perfectly, and because you have y in one and -y in the other, just use the addition method and add the equations. The y's will cancel out, and you'll end up with a single equation in x. Then solve for x. That's all you need to do since all they ask for is the x value of the solution.

OpenStudy (fortytherapper):

Like for the first one \[x + y = 3\] Subtracting x from both sides means doing this: \[(x-x)+y=3-x\] x and -x cancel out, giving you \[y = 3-x\]

OpenStudy (anonymous):

Ok!

OpenStudy (fortytherapper):

How would you start the second one?

OpenStudy (anonymous):

Y = 4x?

OpenStudy (mathstudent55):